You own a portfolio that has $1,589 invested in Stock A and $634 invested in Stock B. If the returns on these stocks are 9.89% and 10.67%, respectively, what is the return on the portfolio?
Answer #1

Total amount invested=1,589+634=2223

Portfolio return=Respective return*Respective weight

=(1,589/2223*9.89)+(634/2223*10.67)

=10.11%(Approx).